The roots of the equation ax squared plus 4x equals -2 are real, rational and equal when a has a value of:
1
Expert's answer
2012-11-16T06:18:56-0500
ax² + 4x + 2 = 0

D = 4² - 4*a*2 = 16-8a

The roots of this equation are real, rational and equal only if discriminant D is equal to zero:

D = 0 <==> 16-8a = 0 <==> a = 2.

So, the roots are real, rational and equal when a has a value of 2.